# Knowledge Independence Breeds Disruption but Limits Recognition

### v1

@@ Abstract
The most well-known research on disruption stems from team science. Prior studies have highlighted the impact of team composition on disruption, emphasizing factors such as team size  (?), geographic distance  (?), and collaboration freshness  (?). These studies suggest that small, onsite, and fresh teams are more likely to produce disruptive research. Based on these insights and our discovery of the link between KI and disruption, we investigate whether such teams tend to build on independent knowledge. Indeed, Fig. 3A-C confirms that this has been the case for decades, indicating that KI may serve as a mediator in producing disruptive research. To explore this possibility, we analyze the relationship between disruption and team composition while controlling for KI. Our investigations (Fig. 3D-F) reveal that teams favoring higher KI systematically produce more disruptive research. However, once KI is controlled for, the slope of the disruption–team relationship becomes notably weaker compared to the uncontrolled case (inset, Fig. 3D-F), verifying the existence of the mediating role of KI. Mediation analyses further confirm the non-trivial mediating effect from KI, particularly when compared to other variables (table S2-S4). These findings suggest that a team’s preference for independent knowledge amplifies its disruptive potential, regardless of its size, geographic distance, or collaboration freshness.
@@
A well-documented phenomenon in the study of scientific impact is the negative correlation between a paper’s disruption index and its citation-based impact  (?, ?, ?). We verify that this pattern persists across journal ranks, publication years, and disciplines (Fig. 4A-C). However, the mechanism underlying this correlation remains to be explored. Since both the disruption and citation are the collective measure of the peer recognition, their correlation is hard to be untangled without linking them to the intrinsic property of the paper itself.
@@
Given the advantages of interconnected knowledge in facilitating knowledge dissemination and adoption  (?, ?), it is reasonable to hypothesize that higher KI may hinder citation-based impact. To test this, for any given year $T$, we measure how KI is distributed across all papers published in that year, thereby forming a baseline distribution. Then, for each subsequent year, $T+t$ ($t\geq 1$), we examine the bibliographies of all papers published in that year, and extract only the references that were published in year $T$. Finally, we compute the KI distribution of the extracted references—where a paper may be counted multiple times if cited more than once—and compare it against the baseline distribution. The results show a systematic shift towards lower KI values, i.e., papers with negative KI values are consistently overrepresented among cited references compared to the baseline distribution, while those with positive KI values are consistently underrepresented (Fig. 4D). This pattern supports our hypothesis, indicating a negative correlation between KI and citation-based impact (Fig. 4E). In addition, as cited age $t$ increases, the KI distribution of yearly cited papers gradually converges towards the baseline distribution, suggesting that the citation disadvantage of higher-KI papers diminishes over time. This delayed recognition indicates that higher-KI papers are more likely to be “sleeping beauties”  (?), remaining dormant for longer periods before experiencing citation bursts (Fig. 4F).
@@
The above analyses confirm that KI is associated with both limited and delayed academic recognition, suggesting that the negative relationship between disruption and impact may be driven by KI. To explore this possibility, we analyze the relationship between disruption and impact while controlling for KI. Our results (Fig. 4G) show that among papers with comparable impact, disruption systematically increases with KI. Notably, once KI is controlled for, the negative slope of the disruption-impact relationship becomes significantly weaker compared to the uncontrolled case (inset, Fig.4G), highlighting KI’s mediating role. Mediation analyses further confirm that KI accounts for a substantial portion of the observed negative relationship, exceeding the explanatory power of other covariates (table S5). These findings confirm our hypothesis that KI increases the challenge of knowledge diffusion while simultaneously enhancing the disruptive potential of scientific work.
@@ Discussion
This paper offers new insights into a perplexing question in scientific innovation: Why has the explosion of knowledge been accompanied by a paradoxical decline  (?) in scientific innovation? Our analysis reveals a steady decline in research that builds on independent knowledge. Yet, such research occupies a favorable niche within the intellectual ecosystem, leveraging latent connections in prior knowledge to produce more disruptive outputs. But this comes at a cost: its overall impact remains limited compared to mainstream research that builds on interconnected knowledge. These findings reveal a universal law in scientific recognition: Knowledge independence breeds disruption but limits recognition.
Multiple studies have examined the relationship between the scientific innovation of a given paper and the topological structure of its bibliography. These studies suggest that the novelty index of the bibliography correlate with the paper’s disruption and impact  (?, ?). While our knowledge independence (KI) index also focuses on the paper’s bibliography, it differs substantially from the aforementioned index. More specifically, the novelty index assesses the atypicality of journal co-appearances within the bibliography, offering a coarse-grained characterization of meso-level relationships between combined knowledge. In contrast, our KI focuses on micro-level relationships by examining direct citation links between references, providing a more granular characterization of knowledge interconnection. The fundamental difference between these measures is reflected in their inverted U-shaped relationship (figure S4A).
@@
Our work also complements prior studies examining the relationship between team composition and scientific disruption, which highlight the disruptive potential of small  (?), onsite  (?), and fresh teams  (?). In particular, we find that such teams exhibit a stronger preference to build on independent knowledge. Moreover, we demonstrate that the relationship between team composition and disruption weakens when controlling for KI. These findings reinforce our assertion that KI serves as a mediator, offering an additional boost to disruption. Consequently, our work suggests a practical strategy for scientists aiming to achieve disruptive discoveries: researchers should prioritize ideas composed of independent knowledge in their collaborative efforts, regardless of team size, location, or collaboration experience.
@@
Finally, our work extends previous observations  (?, ?, ?) on the negative relationship between disruption and impact, demonstrating that this pattern persists across various attributes of papers, including journal rank, publication year, and discipline. But the negative relationship between disruption and impact diminishes when controlling for KI, highlighting the mediating role of knowledge independence in the paradoxical negative relationship between disruption and impact. A deeper insight into the dual role of knowledge independence in scientific innovation is that when a paper is grounded in independent knowledge, the non-redundant information it introduces allows it to function as a knowledge broker, attracting dedicated niche attention. However, the lack of explicit connections between the recombined knowledge simultaneously impedes its broader diffusion.
Although research agendas leveraging independent knowledge can foster scientific disruption and increase the likelihood of receiving prestigious awards (figure S22A), knowledge independence has notably declined across disciplines—from the natural to the social sciences—over the past six decades (Fig. 1C). One possible explanation behind this phenomenon could be the increasing pressure to “publish or perish”  (?, ?, ?). For example, career advancement—through promotions or funding—is often contingent on a steady publication record, which may disincentivize risk-taking in favor of safer, incremental research. We confirm this by observing that funding sources are more dedicated towards papers built on highly dependent knowledge (figure S22B). This funding-driven preference may create a reinforcing cycle that prioritizes well-established knowledge, making it more challenging for researchers to pursue disruptive ideas. Another possible explanation could be the increasing knowledge burden  (?). As the volume of scientific knowledge grows, scholars are increasingly driven to specialize  (?), which may limit their ability to traverse knowledge boundaries effectively  (?). To counteract these barriers, policymakers should consider implementing more flexible evaluation criteria that reward innovative, high-risk research beyond incremental contributions. Additionally, research service providers should develop advanced paper recommendation systems to help scholars discover and integrate independent knowledge more efficiently. Together, these initiatives can encourage research agendas that prioritize independent knowledge, ultimately fostering greater scientific disruption.
Figure 1: Quantifying knowledge independence. (A) Three examples of how knowledge independence (KI) is calculated for a given focal paper (blue diamond) based on its references (gray circles). A reference is of $i$-type (green) if it does not cite any other reference, or of $j$-type (orange) if it does. KI is then the difference in the fraction of $i$-type and $j$-type references. (B) The KI distribution of 29,267,547 papers with at least two references published between 1955 and 2017 in the Web of Science. (C) The evolving downtrend of KI over time across disciplines. Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure 2: Knowledge independence is associated with scientific disruption. (A) For the 27,769,300 papers in the Web of Science that are cited at least once, the average disruption percentile (blue curve, left y-axis) and the disruption positive ratio (green curve, right y-axis) increase with the KI percentile. (B) The association between KI and disruption persists, regardless of the impact of the focal paper. Here, impact is measured as the number of citations received within the first five years after publication, denoted by $C_{5}$. (C) The association between KI and disruption persists across decades. (D) The association between KI and disruption persists, regardless of discipline. Bootstrapped 95% confidence intervals are shown as shaded bands in panels A-D. (E) The ATT (average treatment effect on the treated) matrix of KI on disruption via propensity score matching (PSM). Each controlled group is set as a baseline, and ATTs are calculated for comparisons between the baseline and each of the treated groups. Blue cells represent negative ATTs, while red ones represent positive ATTs, with color intensity proportional to the absolute value. Within each controlled group, the ATT increases as the treated size grows, transitioning from negative to positive when the treated size approaches the controlled size. Each ATT is tested against the null hypothesis (ATT equals 0) using a two-sided t-test. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Figure 3: Team’s preference for knowledge independence. (A-C), The KI percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Curves with larger bounds are displayed in the inset to improve visualization. Over time, the relative KI percentile decreases with team size (panel A) and team distance (panel B), and increases with team freshness (panel C); the only exception is observed for the 1950s in the team distance analysis (inset of panel B). (D-F), Solid lines depict the relationships between disruption and team properties with fixed KI values, alongside the uncontrolled cases (insets). Dashed lines represent linear fitted curves. Teams with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are flatter compared to the uncontrolled curves (insets). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure 4: Knowledge independence explains the negative relationship between disruption and impact. We observe a universal negative relationship between papers’ disruption and impact, regardless of the journal rank (panel A), time period (panel B), and discipline (panel C). Here, a paper’s impact is measured by the percentile of its total citations within the same publication year and discipline, and the journal rank is based on Scimago Journal Rank. (D) For papers published in the year $T=2000$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline (see figure S11 for different values of $T$). (E) The citation percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Over time, the relative citation percentile decreases with the KI percentile. (F) A paper with a higher KI is more likely to become a Sleeping Beauty (SB coefficient, blue curve, left y-axis) and takes a longer silence to usher in the citation burst (Awaking time, green curve, right y-axis). The SB coefficient and Awaking time  (?) capture a delayed scientific recognition by quantifying the convexity and the age of abrupt change in the distribution of annual citations of a paper. (G) Solid lines depict the relationships between disruption and impact with fixed KI values, alongside the uncontrolled case (inset). Dashed lines represent linear fitted curves. For papers of comparable impact, those with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are significantly flatter compared to the uncontrolled curve (inset). Bootstrapped 95% confidence intervals are shown as shaded bands.
Table 1: Normalized regression models of disruption and the influencing factors. We perform normalized ordinary-least-squares (OLS) regression analyses to examine the effect of KI on disruption, while controlling for various confounding factors. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6 Model 7 Model 8 Model 9 Model 10 Model 11
KI 0.4333*** 0.4217*** 0.4332*** 0.4208*** 0.4312*** 0.4328*** 0.2910*** 0.4332*** 0.4317*** 0.4333*** 0.2768***
(0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000)
Log $C_{5}$0.0433*** -0.0318***
(0.000) (0.000)
Novelty 0.0057*** -0.0035
(0.000) (0.000)
Reference count -$0**$0205***
(0.000) (0.000)
Reference age 0.0577*** 0.0094***
(0.000) (0.000)
Reference $C_{5}$0.0205*** -0.0173***
(0.000) (0.000)
Reference disruption 0.2318*** 0.2265***
(0.000) (0.000)
Team size -$0**$0030***
(0.000) (0.000)
Team distance -0.0318*** -0.0231***
(0.000) (0.000)
Team freshness -$0**$0008**
(0.000) (0.000)
Fixed discipline Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Fixed year Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
N 9857467 9857467 9857467 9857467 9857467 9857467 9857467 9857467 9857467 9857467 9857467
$R^{2}$ 0.223 0.224 0.223 0.224 0.226 0.223 0.247 0.223 0.224 0.223 0.249
Acknowledgments
Funding:
The work is supported by National Natural Science Foundation of China (NSFC) (No.72374173), University Innovation Research Group of Chongqing (No.CXQT21005), Fundamental Research Funds for the Central Universities (No.SWU-XDJH202303), China Scholarship Council (CSC) program (No.202306990091), and Chongqing Graduate Research Innovation Project (No.CYB22129).
Author contributions:
X.-Y. Y., T. R. and T. J. conceived and designed the research. X.-Y. Y. analyzed the empirical data and performed the visualizations. X.-Y. Y. and T. J. validated the methods. X.-Y. Y., T. R. and T. J. wrote and edited the manuscript.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
Our study draws on data from the Web of Science (WOS) and SciSciNet. The WOS dataset is not publicly available, and was used under license from their respective publishers. The SciSciNet dataset is accessible via the research  (?).
Supplementary materials
Materials and Methods
Figs. S1 to S22
Tables  S1 to S11
Supplementary Materials for
Knowledge Independence Breeds Disruption but Limits Recognition
Xiaoyao Yu, Talal Rahwan, Tao Jia
This PDF file includes:
Materials and Methods
Figures  S1 to S22
Tables  S1 to S11
Materials and Methods
@@ Acknowledgments
Our data source is from the Web of Science (WOS), encompassing records from the Science Citation Index Expanded (SCIE), Social Sciences Citation Index (SSCI), and Arts & Humanities Citation Index (AHCI) databases. This dataset includes over 53.8 million publications from 1955 to 2017, with 34,353,967 articles, 1,968,358 letters, 1,934,079 proceedings papers, and 1,473,705 reviews. Other types of records that fall out of these “traditional” forms of articles are excluded from our analyses. To classify papers by academic discipline, we adopt the categorization scheme developed by Wu et al.  (?), which consolidates 258 subject categories into 14 major disciplines: Agriculture, Biology, Chemistry, Computer science, Medicine, Engineering, Environment sciences, Humanities, Law, Management, Mathematics, Multidiscipline, Physical sciences, and Social sciences. We also replicate our main findings on an additional dataset, SciSciNet  (?), which is developed from the Microsoft Academic Graph (MAG) dataset  (?). This dataset comprises approximately 134 million publications from 1950 to 2022, thereby complementing the WOS data with a larger data scale and more recent records. Similarly, we consider the “traditional” forms of articles in our analyses, such as journal and conference papers. The results based on SciSciNet are provided in the Supplementary Materials (figure S12-S22, table S6-S11).
@@
Four alternative definitions of knowledge independence are proposed for comparison. KI is the principal definition used in the main text. $\text{KI}^{{}^{\prime}}$ simplifies the equation by only measuring the fraction of $i$-type references. $\text{KI}_{\rm{adj}}$ subtracts 1 from the count of $i$-type references, to account for the oldest reference in the bibliography, which cannot cite any other reference in that list. Finally, $\text{KI}^{{}^{\prime}}_{\rm{adj}}$ simplifies the previous equation by only measuring the fraction of $i$-type reference. More formally, the four alternative measures are defined as follows:
  • $\text{KI}=\dfrac{n_{i}-n_{j}}{n_{i}+n_{j}}$
  • $\text{KI}^{{}^{\prime}}=\dfrac{n_{i}}{n_{i}+n_{j}}$
  • $\text{KI}_{\rm{adj}}=\dfrac{(n_{i}-1)-n_{j}}{(n_{i}-1)+n_{j}}$
  • $\text{KI}^{{}^{\prime}}_{\rm{adj}}=\dfrac{n_{i}-1}{(n_{i}-1)+n_{j}}$
@@
The groundbreaking work of Funk and Owen-Smith  (?) proposed an index that analyzes patent citations to quantify technological change. This index was then advocated by the highly influential work of Wu et al.  (?), who coined the term “disruption index” and were the first to apply it in the analysis of scientific papers. The basic idea behind this index is that a disruptive work shifts the focus of subsequent research from its prior knowledge to itself. Therefore, the disruption index is calculated as $D=\dfrac{c_{i}-c_{j}}{c_{i}+c_{j}+c_{k}}$, where $c_{i}$ represents the number of works that only cite the focal work, $c_{j}$ represents the number of works that cite both the focal work and its references, and $c_{k}$ represents the number of subsequent works that cite the references of the focal work but not the focal work itself. There has been ongoing debate on the rationale behind the inclusion of the term $c_{k}$ in the formula  (?, ?, ?). Specifically, $c_{k}$ quantifies the attention directed towards the focal work’s references while bypassing the focal work itself. Under fixed $c_{i}$ and $c_{j}$, focal works with higher $c_{k}$ values would be considered less disruptive. However, a contradiction arises in cases of negative $D$, where an increase in $c_{k}$ paradoxically increases $D$ rather than decreasing it—contrary to the conceptual definition of disruption. To address this inconsistency, we adopt a modified definition of disruption that excludes the term $c_{k}$. This modified measure serves as the principal metric of disruption in our study, while the original definition is retained as an alternative for robustness check. Additionally, we implement an open time window to count the citations a focal work receives ($c^{o}$) and use a 5-year time window ($c^{5}$) as an alternative measure.

### v2

@@ Results
Figure 3: $|$ Team’s preference for knowledge independence. (A-C) The KI percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Curves with larger bounds are displayed in the inset to improve visualization. Over time, the relative KI percentile decreases with team size (panel A) and team distance (panel B), and increases with team freshness (panel C). Here we calculate the geographic team distance by the coordinate information of affiliations, and categorize teams into two types—onsite ($\leq 100$ km) and remote (>100>100 km) teams (panel B). The relationship between KI and average collaboration distance is displayed in inset of panel B. (D-F) Solid lines depict the relationships between disruption and team properties with fixed KI values, alongside the uncontrolled cases (insets). Dashed lines represent linear fitted curves. Teams with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are flatter compared to the uncontrolled curves (insets). Bootstrapped 95% confidence intervals are shown as shaded bands.
The most well-known research on disruption stems from team science. Prior studies have highlighted the impact of team composition on disruption, emphasizing factors such as team size  [wu2019large], geographic distance  [lin2023remote], and collaboration freshness  [zeng2021fresh]. These studies suggest that small, onsite, and fresh teams are more likely to produce disruptive research. Based on these insights and our discovery of the link between KI and disruption, we investigate whether such teams tend to build on independent knowledge. Indeed, Fig. 3A-C confirms that this has been the case for decades, indicating that KI may serve as a mediator in producing disruptive research. To explore this possibility, we analyze the relationship between disruption and team composition while controlling for KI. Our investigations (Fig. 3D-F) reveal that teams favoring higher KI systematically produce more disruptive research. However, once KI is controlled for, the slope of the disruption–team relationship becomes notably weaker compared to the uncontrolled case (inset, Fig. 3D-F), indicating the existence of the mediating role of KI. Mediation analyses  [nguyen2021clarifying] (details in Methods) further confirm the non-trivial mediating effect from KI, particularly when compared to other variables (Extended Data Table 1-3). These findings persist when applied to alternative disruption metrics (Table S2-S4 in Section S5), suggesting that a team’s preference for independent knowledge amplifies its disruptive potential, regardless of its size, geographic distance, or collaboration freshness.
@@
A well-documented phenomenon in the study of scientific impact is the negative correlation between a paper’s disruption index and its citation-based impact  [bornmann2020disruption, liang2023bias, zeng2023disruptive]. We provide evidence indicating that this pattern persists across journal ranks, publication years, and disciplines (Fig. 4A-C). However, the mechanism underlying this correlation remains to be explored. Considering that both the disruption and citation are the extrinsic measures of the collective peer recognition, it indicates that there might be some symbiosis among these citation-based extrinsic metrics, and the correlation between them is likely to be a spurious or biased association stemming from a more intrinsic attribute.
Figure 4: $|$ Knowledge independence explains the negative relationship between disruption and impact. We observe a universal negative relationship between papers’ disruption and impact, regardless of the journal rank (panel A), time period (panel B), and discipline (panel C). Here, a paper’s impact is measured by the percentile of its total citations within the same publication year and discipline, and the journal rank is based on Scimago Journal Rank. (D) For papers published in the year $T=2000$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline (see Extended Data Fig. 12 for different values of $T$). (E) The citation percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Over time, the relative citation percentile decreases with the KI percentile. (F) A paper with a higher KI is more likely to become a Sleeping Beauty (SB coefficient, blue curve, left y-axis) and takes a longer silence to usher in the citation burst (Awaking time, green curve, right y-axis). The SB coefficient and Awaking time  [ke2015defining] capture a delayed scientific recognition by quantifying the convexity and the age of abrupt change in the distribution of annual citations of a paper. (G) Solid lines depict the relationships between disruption and impact with fixed KI values, alongside the uncontrolled case (inset). Dashed lines represent linear fitted curves. For papers of comparable impact, those with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are significantly flatter compared to the uncontrolled curve (inset). Bootstrapped 95% confidence intervals are shown as shaded bands.
@@
Given the advantages of interconnected knowledge in facilitating knowledge dissemination and adoption  [nooteboom2009cognitive, deichmann2020ideas], it is reasonable to hypothesize that higher KI may hinder citation-based impact. To test this, for any given year $T$, we measure how KI is distributed across all papers published in that year, thereby forming a baseline distribution. Then, for each subsequent year, $T+t$ ($t\geq 1$), we examine the bibliographies of all papers published in that year, and extract only the references that were published in year $T$. Finally, we compute the KI distribution of the extracted references—where a paper may be counted multiple times if cited more than once—and compare it against the baseline distribution. The results show a systematic shift towards lower KI values, i.e., papers with negative KI values are consistently overrepresented among cited references compared to the baseline distribution, while those with positive KI values are consistently underrepresented (Fig. 4D). This pattern supports our hypothesis, indicating a negative correlation between KI and citation-based impact (Fig. 4E). In addition, as cited age $t$ increases, the KI distribution of yearly cited papers gradually converges towards the baseline distribution, suggesting that the citation disadvantage of higher-KI papers diminishes over time. This delayed recognition indicates that higher-KI papers are more likely to be “sleeping beauties”  [ke2015defining], remaining dormant for longer periods before experiencing citation bursts (Fig. 4F).
@@
The analyses above verify that KI is associated with both limited and delayed academic recognition, suggesting that the negative relationship between disruption and impact may be influenced by KI. Given that both disruption and recognition are ex-post variables relative to KI, their relationship is more potentially and logically depicted as being confounded by KI, an ex-ante variable that precedes and affects both outcomes. To explore this possibility, we analyze the relationship between disruption and impact while controlling for KI. Our results (Fig. 4G) show that among papers with comparable impact, disruption systematically increases with KI. Notably, once KI is controlled for, the negative slope of the disruption-impact relationship becomes significantly weaker compared to the uncontrolled case (inset, Fig. 4G), highlighting the confounding potential of KI. Further confounding analyses reveal that KI accounts for a substantial portion of this observed negative relationship, surpassing the explanatory power of other factors (Extended Data Table 4), which provides additional analytical evidence for KI’s confounding role. These findings persist when applied to alternative metrics of disruption (Table S5 in Section S5), supporting our hypothesis that KI increases the challenge of knowledge diffusion while simultaneously enhancing the disruptive potential of scientific work.
@@ Discussion
To explain why disruption declined systematically and associated with fewer citations and delayed recognition, we identify an intrinsic property of scientific papers, knowledge independence (KI), that reliably predicts disruption. We also show that this property mediates the disruptive advantage of small, onsite, and fresh teams, confirming its role as a core driver of disruptive processes. Moreover, the decades-long decline of KI provides a mechanistic explanation for the parallel decline in disruption. Finally, causal inference methods coupled with Monte Carlo simulations, suggest that KI is the underlying driver behind the persistent trade-off between disruption and impact.
@@
Our work complements prior studies examining the relationship between team composition and scientific disruption, which highlight the disruptive potential of small  [wu2019large], onsite  [lin2023remote], and fresh teams  [zeng2021fresh]. In particular, we find that such teams exhibit a stronger preference to build on independent knowledge. Moreover, we demonstrate that the relationship between team composition and disruption weakens when controlling for KI. These findings highlight the role KI as a mediator, offering an additional boost to disruption. Consequently, our work suggests a practical strategy for scientists aiming to achieve disruptive discoveries: researchers should prioritize ideas composed of previously unlinked but implicitly related knowledge in their collaborative efforts, regardless of team size, location, or collaboration experience.
The decades-long decline of KI provides a mechanistic explanation for the decline in disruption. One may argue that this parallel decline might be caused by mere chance, i.e., by the structural evolution of the citation network itself. We rule out this possibility with our randomized rewiring Monte Carlo analysis (see Section S2.1 in Supplementary Materials): the randomized network fails to replicate the decline of KI, signaling that it is a consequence of a behavioural shift in the way knowledge is combined. The randomized network also fails to reproduce the KI-disruption association, further indicating that the phenomenon we observe is derived from non-trivial scientists’ behaviour. This temporal alignment—where the intrinsic property (KI) and the extrinsic outcome (disruption) decline in parallel—provides a plausible explanation for why disruption is decreasing.
@@
Moreover, our work extends previous observations  [bornmann2020disruption, liang2023bias, zeng2023disruptive] on the negative relationship between disruption and impact, demonstrating that this pattern persists across journal rank, publication year, and discipline. More importantly, we offer a novel mechanistic explanation for this phenomenon. Our analysis reveals that the seemingly negative relationship between disruption and impact is confounded by an intrinsic property of the paper: its knowledge independence (KI). When controlling for KI, this relationship is substantially attenuated. Further confounding analysis confirms that KI plays a dominant role, suggesting that it drives the observed relationship between these two extrinsic, citation-based metrics. Intuitively, KI plays a dual role: when a paper is grounded in independent knowledge, the non-redundant information it introduces allows it to function as a “knowledge broker”, attracting dedicated niche attention and thus increasing its disruptive potential. However, the very lack of explicit connections between the combined knowledge simultaneously impedes its broader diffusion, leading to lower and more delayed recognition.
To rigorously test our mechanism against the null hypothesis of trivial structural determinism—the idea that our finding is a mere artifact of random network evolution—we conducted a suite of Monte Carlo simulations. First, our falsification test (see Section S2.1 in Supplementary Materials), which randomizes the empirical network while preserving its core structure, finds that the KI-Disruption and KI-Impact correlations completely disappear. This provides evidence that our findings are not driven by a trivial, universal network property but a substantive, but rather by non-random pattern of scientific behavior. Second, our generative Monte Carlo simulations (see Section S2.2 in Supplementary Materials) serves to validate the explanatory sufficiency of our proposed mechanism, rather than to identify predictors. We find that standard network models cannot reproduce this phenomenon. By embedding the specific cognitive preference for independent knowledge—which was empirically isolated as the dominant driver in our earlier analyses—into the model, we demonstrate that this mechanism is sufficient to reproduce the work that is highly disruptive but less recognized. This comprehensive evidence, from large-scale statistical observations to targeted mechanistic simulations, supports our conclusion that a paper’s intrinsic knowledge independence is the non-trivial, human-driven missing link that explains the trade-off between disruption and recognition in science.
Methods
@@
Our data source is from the OpenAlex  [priem2022openalex], which includes over 114 million publications from 1950 to 2024, with 92,659,091 articles, 3,411,018 reviews, 2,693,331 preprints, and 1,391,755 letters. Other types of records that fall out of these “traditional” forms of articles are excluded from our analyses. The OpenAlex consolidates 244 subject categories into 26 major fields of discipline: Energy, Physics, Biology, Nursing, Business, Medicine, Dentistry, Economics, Chemistry, Psychology, Veterinary, Immunology, Mathematics, Engineering, Agriculture, Neuroscience, Pharmacology, Earth Sciences, Social Sciences, Computer Science, Materials Science, Decision Sciences, Health Professions, Arts and Humanities, Chemical Engineering, and Environmental Science. We also replicate our main findings on two additional datasets, SciSciNet  [lin2023sciscinet] and Web of Science (WOS). The SciSciNet dataset comprises approximately 130 million publications from 1950 to 2022, and the WOS dataset comprises approximately 53.8 million publications from 1955 to 2017. Similarly, we consider the “traditional” forms of articles in our analyses, such as journal and conference papers. The results based on SciSciNet and WOS datasets are provided in the Supplementary Materials (Section S6, S7). We further extend our analysis on the relationship between KI and disruption to technical area based on OECD patent dataset (latest version: Feb-2025), which comprises approximately 30 million patents filed at the European Patent Office (EPO) and the United States Patent and Trademark Office (USPTO), or via the Patent Co-operation Treaty (PCT) from 1978 to 2024. The results based on OECD patent dataset are provided in the Supplementary Materials (Section S8).
@@
Four alternative measures of knowledge independence are proposed for comparison. KI is the principal definition used in the main text. $\text{KI}^{{}^{\prime}}$ simplifies the equation by only measuring the fraction of indind-type references. $\text{KI}_{\rm{adj}}$ subtracts 1 from the count of indind-type references, to account for the oldest reference in the bibliography, which cannot cite any other reference in that list. Finally, $\text{KI}^{{}^{\prime}}_{\rm{adj}}$ simplifies the previous equation by only measuring the fraction of indind-type reference. More formally, the four alternative measures are defined as follows:
  • $\text{KI}=\dfrac{n_{\rm ind}-n_{\rm dep}}{n_{\rm ind}+n_{\rm dep}}$
  • $\text{KI}^{{}^{\prime}}=\dfrac{n_{\rm ind}}{n_{\rm ind}+n_{\rm dep}}$
  • $\text{KI}_{\rm{adj}}=\dfrac{(n_{i}-1)-n_{\rm dep}}{(n_{i}-1)+n_{\rm dep}}$
  • $\text{KI}^{{}^{\prime}}_{\rm{adj}}=\dfrac{n_{\rm ind}-1}{(n_{i}-1)+n_{\rm dep}}$
@@
The groundbreaking work of Funk and Owen-Smith  [funk2017dynamic] proposed an index that analyzes patent citations to quantify technological change. This index was then advocated by the highly influential work of Wu et al.  [wu2019large], who coined the term “disruption index” and were the first to apply it in the analysis of scientific papers. The basic idea behind this index is that a disruptive work shifts the focus of subsequent research from its prior knowledge to itself. Therefore, the disruption index is calculated as $D=\dfrac{c_{i}-c_{j}}{c_{i}+c_{j}+c_{k}}$, where $c_{i}$ represents the number of works that only cite the focal work, $c_{j}$ represents the number of works that cite both the focal work and its references, and $c_{k}$ represents the number of subsequent works that cite the references of the focal work but not the focal work itself. There has been ongoing debate on the rationale behind the inclusion of the term $c_{k}$ in the formula  [wu2019confusing, bornmann2020disruption, leibel2024we]. Specifically, $c_{k}$ quantifies the attention directed towards the focal work’s references while bypassing the focal work itself. Under fixed $c_{i}$ and $c_{j}$, focal works with higher $c_{k}$ values would be considered less disruptive. However, a contradiction arises in cases of negative $D$, where an increase in $c_{k}$ paradoxically increases $D$ rather than decreasing it—contrary to the conceptual definition of disruption. To address this inconsistency, we adopt a modified definition of disruption that excludes the term $c_{k}$. This modified measure serves as the principal metric of disruption in our study, while the original definition is retained as an alternative for robustness check. Additionally, we implement an open time window to count the citations a focal work receives ($c^{o}$) and use a 5-year time window ($c^{5}$) as an alternative measure.